An automatic coffee machine is programmed so that it produced an average of 200 ml per cup. The amount of coffee follows a normal distribution with standard deviation equal to 10ml. if 25% of the drinks are below k value, find the value of k?



1
Expert's answer
2021-01-21T00:20:30-0500

"P(Z<z)=0.25\to z=-0.6745."

"\frac{k-\mu}{\sigma}=z."

"\frac{k-200}{10}=-0.6745."

"k=200-0.6745*10=193.255\;ml."